A battle has erupted over a local power plant owner's decision to raise the power delivered by the plant from 255 MW to 726 MW (The delivered power accounts for energy losses). The local population is increasing, so the area needs more power. However, the plant expends its heated waste water* into the Frustration River; local citizens are concerned that the heated water will raise water temperatures to the point where the endangered river smelt will die from oxygen depletion. Local biologists called into the fray report that the smelt can only withstand an average increase of 3.44°C. The river is 212 ft wide at the point where the waste water is dumped, but the depth and cross section is not well known. Farther downstream, however, the water flows through a narrow man-made concrete channel of rectangular cross-section, which has a width of 84.7 ft and depth of 43.2 ft. The water flows through this channel at 12.4 mph. The power plant claims a planned efficiency of 37.30% for its new power plant design. Based on this information, what is the maximum power that the power plant can output without endangering the river smelt?
Answer #1

given
initial power output, Pi = 255 MW
Pf = 726 MW

width of river, w = 212 ft

dimensions of channel axb = 84.7 ft x 43.2 ft = 3659.04 ft^2 = 339.936 m^2
v = 12.4 mph = 5.588 m/s
hence mass flow rate = rho*axb*v
rho = 1000 kg/m^3 ( density of water)
m' = 1899562.368 kg/s

efficiency = 37.3 %
power output = P
Wasted power = Pw
then
e = P/(P + Pw)
1/e = 1 + Pw/P
Pw = (1/e - 1)*P

now Pw = m'cdT
c = 4186 J/kg K ( heat capacity of water)
dT = 3.44 K or C
hence
Pw = 27353394169.22112 W = 27353.39 MW
P = 45980102263.007 W = 45980 MW

Power transmitted = P -Pw = 18626 MW
hence the plant can increase power output to 18626+255 = 18881 MW without endangering the local species in water